OpenStudy (katnissstyles000):

HELP!!!!!!!WILL FAN AND MEDAL!!!! 7e−4=31

OpenStudy (amina_soneviseth):

Simplifying 7e + -4 = 31 Reorder the terms: -4 + 7e = 31 Solving -4 + 7e = 31 Solving for variable 'e'. Move all terms containing e to the left, all other terms to the right. Add '4' to each side of the equation. -4 + 4 + 7e = 31 + 4 Combine like terms: -4 + 4 = 0 0 + 7e = 31 + 4 7e = 31 + 4 Combine like terms: 31 + 4 = 35 7e = 35 Divide each side by '7'. e = 5 Simplifying e = 5

OpenStudy (anonymous):

E= 5

OpenStudy (katnissstyles000):

please help!!!!

OpenStudy (amina_soneviseth):

@KatnissStyles000 does that help? would love a fan and medal!

OpenStudy (katnissstyles000):

thank you

OpenStudy (anonymous):

Welcome

OpenStudy (amina_soneviseth):

@KatnissStyles000 no problem! :D

OpenStudy (katnissstyles000):

love yall ur livesavers!!!!